THEORETICAL SPIN-ORBIT SPECTROSCOPY OF Cl DOPANTS IN SOLID PARAHYDROGEN

\underline{ROBERT HINDE}, Department of Chemistry, University of Tennessee, Knoxville, TN 37996 USA.

Solid parahydrogen (pH_{2}) matrices containing open-shell (^{2}P) Cl and Br atoms as substitutional impurities exhibit several infrared (IR) absorption features associated with intermolecular interactions between the halogen atom and nearby pH_{2} molecules.^{a} These dopant-induced IR absorption features are associated with (i) spin-orbit (SO) excitation of the halogen atom, and (ii) cooperative excitations in which a single IR photon both triggers SO excitation of the halogen atom and excites the H-H stretching coordinate of a nearby pH_{2} molecule. The IR spectral features thus contain detailed information about the interaction between the open-shell halogen atom and the surrounding pH_{2} molecules, and in particular tell us about the coupling between the atom’s electronic degrees of freedom and the nuclear motion of its H_{2} neighbors.

Here we present recent progress towards understanding these features in Cl-doped solid pH_{2}. This system represents an arena for testing theoretical approaches that generate many-body, non-pairwise-additive potential energy functions for the dopant–matrix interaction by combining multiple low-lying Cl–H_{2} pair potential energy functions. Our analysis must account for the fact that the individual pH_{2} molecules in the doped solid undergo large amplitude zero-point motions;^{b} we use quantum Monte Carlo simulations^{c} to model these zero-point motions. We interpret our findings in terms of a reduced dimensionality model that provides a qualitative understanding of the role that matrix zero-point motions play in the observed spectra.

aP. L. Raston and D. T. Anderson, J. Chem. Phys. 126, 021106 (2007); S. C. Kettwich, L. O. Paulson, P. L. Raston and D. T. Anderson, J. Phys. Chem. A 112, 11153 (2008).

bI .F. Silvera, Rev. Mod. Phys. 52, 393 (1980).

cR. J. Hinde, Comput. Phys. Commun. 182, 2339 (2011).